# A Hybrid Data-Driven Multi-Stage Deep Learning Framework for Enhanced Nuclear Reactor Power Prediction

### v3

@@ 2.2.1 Dataset Normalization
The single observation normalization method is included as 5.6. This method could then be applied to the entire dataset to form the input data of the deep learning model. The equations used in the normalization process are shown below.
@@
If all samples were properly classified, undersampling methods could be used. Undersampling is important for models which might have much more of one output than another. The model could preferentially select the more populated class resulting in underfitting. An undersampling method was applied to the generated dataset such that all output classes would have the same number of samples. The undersampling method is included as 5.7.
@@ 2.3.1 Neural Network Design
The data input to some models was different than others. Some models used control rod heights as inputs, while some used total reactivity generated by control rods. Some models reduced a total number of input variables by summing the control rod information (when relevant) into a single variable. If the control rod heights were used in this way, only control rods 1, 2, and 3 were summed. If the reactivity provided by each control rod was used, all of the rods were summed into a single variable. Table 3 shows the types of models designed and trained.
@@
Regression models were created to make a final prediction. The probabilistic output of the classifier networks could help form a better basis to distribute the output values in a regression model, and as such, they were used as input. These models utilized a sigmoid output function in order to squash the output so it could not produce a value greater than 1, which is the normalized maximum power. The architectures were designed similarly to the classifier models, with some having AIO input while others had split input and internal hidden layers. These models were also designed as feedforward neural networks using backpropagation. In both cases, the 5-label output vector from a classifier was used as an additional input. The two general styles are shown in Appendices 5.3 and 5.4, along with the types of models produced in Table 4. All regression models utilized the power change direction. All regression models utilized a mean absolute error loss function.
@@ 2.3.2 Model Hyperparameters
All models were trained using the ADAM optimizer, with a learning rate of 0.0005. Additionally, L2 regularization was utilized in classification model output layers and L1 regularization was utilized in regression model output layers. L2 regularization was utilized for classification models to avoid reducing connection weights to 0., ensuring the models would still have the capability of making a prediction in any category.
Classification models utilized the softmax output activation, in order to produce an output vector of probabilistic nature. The value of the softmax in a given bin is the predicted effective likelihood that the final reactor power would be in that region. Categorical crossentropy loss metrics were utilized, with validation using categorical accuracy.
Regression models utilized sigmoid output activation. The sigmoid function is highly non-linear and produces values in a range from 0 to 1. Since reactor power was log-normalized and could not be below 0W power, the sigmoid function would always produce values in the relevant region while extracting non-linear behavior from the previous layer. Mean absolute error loss metrics were utilized, with validation using MAPE.
All hidden layers in the models utilized hyperbolic tangent activation. Hyperbolic tangent is a highly non-linear function with values ranging from -1 to 1. This allows for a greater diversification in neuron connection weights and can produce negative correlations in models, such as when making a positive versus negative power change.
Due to the number and nature of the models, the number of neurons is omitted here. Layer units may be seen in the model graphs, included in the appendix. In case of further questions about model architecture and design, it is encouraged to contact the authors.
@@ 3 DATA AND RESULTS
After models were created and inputs were generated, the proposed framework could be used in power prediction. The following sections show the training and testing results, with the results graphed for ease of visualization.
@@ 3.1 Training
The table below shows the models trained and relevant variables associated with their training. Models were stopped early if the aforementioned early stopping criterion was met. If a model were stopped this way, it would revert to the epoch in which it had the highest validation accuracy. Information from these epochs is included in Table 5.
@@
The loss and accuracy metrics of the models trained are included as Figures 4 and 5. A history of the training cycle was retained for this purpose. Figure 4 depicts the training losses of each model per epoch of training, and figure 5 depicts the accuracy of each model per epoch trained. This can be used to determine how quickly and effectively the models are trained. High-performance models were saved in case they would need to be loaded in future instances.
It can be seen that while the split-input type models performed well, the AIO input models trained poorly. This is likely due to a lack of feature separation, causing confusion in the model. This would imply that input separation improves training for a physical type of problem.
@@ 3.2.1 Classifiers
After training, the better performing models could be tested with the actual data from the reactor. Data from the reactor could be formatted into vectors and used by the model to predict outputs. Confusion matrices were used to visualize how well the data performed on the original dataset. The result is shown in Figure 6.
@@ 3.2.2 Regression
The regression models were trained using the output from the classifier most closely resembling its input type. The distribution and error of the regression models can be analyzed to show their efficacy. The results are shown in Figures 7, 8, and 9.
@@ 3.2.3 Combined Model
To complete the model, the most successful models from each stage can be combined. A model can be defined in Tensorflow using a trained model from each stage as a functional layer. The result is a model that takes the input utilized in the first stage of the model and internally processes it for use in the second stage of the model. The output of the first stage is also utilized for the input in the second stage, as outlined in the training section. The full model with two functional layers and an input layer can be seen in Figure 11. The selected first-stage model was a1, and the selected second-stage model was b2.
@@
The resultant DNN outputs both the categorical and numerical predictions. Since DNNs are replicable, the predictions in each stage are the same as the predictions from models a1 and b2. This final DNN can be functionally applied or retrofitted and exported to function as a component in a larger system.
@@ 4 DISCUSSION
Practically, a similar type of model could likely be developed in other reactors. The inputs would have to be developed based on the design of the reactor core, and the inputs could further be expanded as well. The normalization of control rod heights could be the cause of the reduced accuracy of the models using control rod heights as input. More likely, however, is that the models struggle to infer the physics of the reactor core. With the reactivity provided, the model does not have to infer the effect of the rods on the power but instead can directly map the reactor power to reactivity. This has the benefit of allowing future models to account for negative reactivity provided by burnable poisons or testing samples. Furthermore, using reactivity would also allow the model to be more useable on reactor core configurations it was not trained on. That said, classifier accuracy for control rods was not impractically low, and most incorrect predictions were in an adjacent energy resolution. The classification accuracy using control rods is at least a proof of concept for future models using similar inputs.
Over the time the data was collected, MSTR saw significant use. This could have led to changes in the core physics. This time period corresponds to a total burnup of 120MWd/MTU. Based on the fuel designs, this is 13% of the fuel’s lifetime burnup. This physical change, along with the increase of poisons generated inside the fuel, could cause complications with the results, and additional support modules may need to be developed to account for these effects.
@@
The other subject of future research would be continued development. Besides developing different architectures for testing, the logical next steps would be either a full predictor, modular controller, or both. A Long Short-Term Memory (LSTM) or convolutional neural network could be developed as a time series predictor. A state-based predictor like the one developed could be used in conjunction with this type of model, which would provide an end state as a horizontal asymptote for the time series. If this type of model was desired, the state-based model could also be modified to produce a value for the time required to reach the final power as well. If an automatic controller were desired, one could be developed in conjunction with a state-based predictor as a sort of look-ahead function. The controller would perform an action, and the predictor could generate the final power of the transient and provide feedback to the controller. In fact, using a similar model to approximate feedback terms from reactor operation could be used in conjunction with other types of models (machine learning or physics-based) to streamline the prediction process. The predictor controller could then be modularized and actively receive live information from the reactor to form an online automatic controller.
@@ 5 CONCLUSIONS & FUTURE WORKS
The analysis of the generated models reveals that the classifier models, particularly those leveraging reactivity inputs, exhibit exceptional performance. The probabilistic nature of the softmax output function facilitates a nuanced evaluation of the model’s predictions, further enhancing its practical utility. The preference for reactivity-based models over those utilizing control rod heights underscores the significance of incorporating physically interpretable parameters in predictive modeling. While the inclusion of power change direction yielded marginal gains in classifier accuracy, its omission in future iterations can potentially simplify the framework without compromising performance. The regression models also demonstrated commendable accuracy, with predictions based on accurate classifier outputs exhibiting a minimum MAPE of 2.3%. The observed increase in prediction error near the maximum power level, attributed to the logarithmic normalization technique, and the occasional 20% errors in regression models, potentially linked to power state clustering, highlight areas for future refinement. The insights gleaned from this study lay a robust groundwork for the application of deep learning in nuclear reactor modeling and control, promising advancements in operational efficiency and safety.
The continuation of this study will contribute to the development of an intelligent digital twin by integrating advanced deep learning frameworks to enable real-time monitoring and optimization. Future iterations of the proposed framework could incorporate
@@
  • A robust uncertainty quantification (UQ) framework (derived from authors’ previous studies Kobayashi et al. (2024c); Kumar et al. (2019, 2022)) is essential for ensuring reliable predictions. This could involve Monte Carlo simulations or Bayesian techniques to account for uncertainties in both real-world reactor data and the machine learning models.

### v4

@@ 2.2.1 Dataset Normalization
The single observation normalization method is included as 5.7. This method could then be applied to the entire dataset to form the input data of the deep learning model. The equations used in the normalization process are shown below.
@@
If all samples were properly classified, undersampling methods could be used. Undersampling is important for models which might have much more of one output than another. The model could preferentially select the more populated class resulting in underfitting. An undersampling method was applied to the generated dataset such that all output classes would have the same number of samples. The undersampling method is included as 5.8.
@@ 2.3.1 Neural Network Design
The data input to some models was different than others. Some models used control rod heights as inputs, while some used total reactivity generated by control rods. Some models reduced the total number of input variables by summing the control rod information (when relevant) into a single variable. If the control rod heights were used this way, only control rods 1, 2, and 3 were summed. If the reactivity provided by each control rod was used, all of the rods were summed into a single variable. Table 3 shows the types of models designed and trained.
@@
Regression models were created to make a final prediction. The probabilistic output of the classifier networks could help form a better basis to distribute the output values in a regression model, and as such, they were used as input. These models utilized a sigmoid output function to squash the output so it could not produce a value greater than 1, which is the normalized maximum power. The architectures were designed similarly to the classifier models, with some having AIO input while others had split input and internal hidden layers. These models were also designed as feedforward neural networks using backpropagation. In both cases, the 5-label output vector from a classifier was used as an additional input. The two general styles are shown in Appendices 5.3 and 5.4, along with the types of models produced in Table 4. All regression models utilized the power change direction. All regression models utilized a mean absolute error loss function.
@@ 2.3.2 Model Hyperparameters
All models were trained using the ADAM optimizer, with a learning rate of 0.0005. Additionally, L2 regularization was utilized in classification model output layers and L1 regularization was utilized in regression model output layers. L2 regularization was utilized for classification models to avoid reducing connection weights to zero, ensuring the models would still have the capability of making a prediction in any category.
Classification models utilized the softmax output activation to produce a probabilistic output vector. The value of the softmax in a given bin is the predicted effective likelihood that the final reactor power would be in that region. Categorical cross-entropy loss metrics were utilized, with validation using categorical accuracy.
Regression models utilized sigmoid output activation. The sigmoid function is highly non-linear and produces values from 0 to 1. Since reactor power was log-normalized and could not be below 0W power, the sigmoid function would always produce values in the relevant region while extracting non-linear behavior from the previous layer. Mean absolute error loss metrics were utilized, with validation using MAPE.
All hidden layers in the models utilized hyperbolic tangent activation. Hyperbolic tangent is a highly non-linear function with values ranging from -1 to 1. This allows for greater diversification in neuron connection weights and can produce negative correlations in models, such as when making a positive versus negative power change.
Due to the number and nature of the models, the number of neurons is omitted here. Layer units may be seen in the model graphs included in the appendix.
@@ 3 DATA AND RESULTS
After models were created and inputs were generated, the proposed framework could be used in power prediction. As part of a digital twin framework, these models provide synchronized and accurate representations of reactor states, supporting real-time decision-making and operational optimization. The following sections show the training and testing results, with the results graphed for ease of visualization.
@@ 3.1 Training
The table below shows the models trained and relevant variables associated with their training. Models were stopped early if the aforementioned early stopping criterion was met. If a model were stopped this way, it would revert to the epoch with the highest validation accuracy. This ensures the trained models retain their performance and robustness for integration into the digital twin, maintaining reliability during operational use. Table 5 includes information from these epochs.
@@
The loss and accuracy metrics of the models trained are included as Figures 4 and 5. A history of the training cycle was retained for this purpose. Figure 4 depicts the training losses of each model per epoch of training, and figure 5 depicts the accuracy of each model per epoch trained. These results illustrate the ability of digital twin deployable models to rapidly converge and achieve high accuracy, demonstrating their suitability for real-time deployment. High-performance models were saved in case they would need to be loaded in future instances.
It can be seen that while the split-input type models performed well, the AIO input models trained poorly. This is likely due to a lack of feature separation, causing confusion in the model. This observation demonstrates the importance of preserving physical structure in the input data, a key requirement for maintaining a digital twin’s fidelity to the physical reactor system. This would imply that input separation improves training for a physical type of problem.
@@ 3.2.1 Classifiers
After training, the better performing models could be tested with the actual data from the reactor. Data from the reactor could be formatted into vectors and used by the model to predict outputs. A digital twin may leverage these classifier models to identify operational states and transitions, critical for dynamic synchronization with the physical system. Confusion matrices were used to visualize how well the data performed on the original dataset. The result is shown in Figure 6.
@@ 3.2.2 Regression
The regression models were trained using the output from the classifier most closely resembling its input type. These regression models enable the digital twin to predict reactor power dynamics with high accuracy, providing actionable insights for operational decision-making. The distribution and error of the regression models can be analyzed to show their efficacy. The results are shown in Figures 7, 8, and 9.
@@ 3.2.3 Combined Model
To complete the model, the most successful models from each stage can be combined. This integration represents the final step in constructing a digital twin that provides both categorical and continuous predictions of reactor states. A model can be defined in Tensorflow using a trained model from each stage as a functional layer. The result is a model that takes the input utilized in the first stage of the model and internally processes it for use in the second stage of the model. The output of the first stage is also utilized for the input in the second stage, as outlined in the training section. The full model with two functional layers and an input layer can be seen in Figure 11. The selected first-stage model was a1, and the selected second-stage model was b2.
@@
This final model embodies digital twin supporting modules by providing synchronized and accurate predictions of reactor behavior, enabling real-time monitoring, predictive diagnostics, and operational optimization. The resultant DNN outputs both the categorical and numerical predictions. Since DNNs are replicable, the predictions in each stage are the same as the predictions from models a1 and b2. This final DNN can be functionally applied or retrofitted and exported to function as a component in a larger system.
@@ 4 DISCUSSION
Practically, a similar type of model could likely be developed in other reactors. As part of a digital twin framework, these models could serve as adaptable virtual replicas tailored to the unique design of each reactor core. The inputs would have to be developed based on the design of the reactor core, and the inputs could further be expanded as well. The normalization of control rod heights could cause the models’ reduced accuracy using control rod heights as input. More likely, however, is that the model needs more feature representation to infer the physics of the reactor core. In a digital twin, the integration of reactivity as an input eliminates this limitation by directly linking the reactor power to reactivity, thereby enhancing the twin’s ability to mirror reactor behavior accurately. This allows future models to account for negative reactivity provided by burnable poisons or testing samples. Furthermore, the use of reactivity enables the digital twin to generalize across different reactor core configurations, ensuring its applicability beyond the conditions on which it was originally trained. That said, classifier accuracy for control rods was not impractically low, and most incorrect predictions were in an adjacent energy resolution. The classification accuracy using control rods is at least a proof of concept for future models using similar inputs.
Over the time the data was collected, MSTR saw significant use. This could have led to changes in the core physics. This time period corresponds to a total burnup of 120MWd/MTU. Based on the fuel designs, this is 13% of the fuel’s lifetime burnup. These evolving reactor conditions highlight the importance of dynamic adaptability within a digital twin framework. Future iterations of these types of models in digital twin systems could incorporate support modules that account for changes in core physics due to burnup, fuel poisoning, or operational conditions, maintaining accuracy over the reactor’s lifecycle.
@@
The other subject of future research would be continued development. Enhancing these models for further deployment in digital twin systems could involve exploring advanced architectures, such as Long Short-Term Memory (LSTM) networks or convolutional neural networks (CNNs), for dynamic time series prediction. A state-based predictor like the one developed could be used in conjunction with this type of model, which would provide an end state as a horizontal asymptote for the time series. This integration would allow the digital twin to simulate reactor transitions dynamically while also providing an estimated time to reach the final state, further aligning with real-time operational needs. If an automatic controller were desired, one could be developed in conjunction with a state-based predictor as a sort of look-ahead function. Such a controller, integrated with the digital twin, could predict reactor state transitions in real time and provide operational feedback to optimize control actions. The controller would perform an action, and the predictor could generate the final power of the transient and provide feedback to the controller. This modularization would allow the digital twin to operate as an online, automated monitoring and control system, dynamically adapting to live reactor conditions and supporting operational decision-making in real time. In fact, using a similar model to approximate feedback terms from reactor operation could be used in conjunction with other types of models (machine learning or physics-based) to streamline the prediction process. The predictor controller could then be modularized and actively receive live information from the reactor to form an online automatic controller.
@@ 5 CONCLUSIONS & FUTURE WORKS
The analysis of the generated models reveals that the classifier models, particularly those leveraging reactivity inputs, exhibit exceptional performance. These models align closely with the principles of a digital twin by providing real-time, probabilistic insights into reactor state transitions, enabling dynamic synchronization with the physical system. The probabilistic nature of the softmax output function facilitates a nuanced evaluation of the model’s predictions, further enhancing its practical utility. The preference for reactivity-based models over those utilizing control rod heights underscores the importance of integrating physically interpretable parameters, a key requirement for ensuring the fidelity of digital twin representations. While the inclusion of power change direction yielded marginal gains in classifier accuracy, its omission in future iterations can potentially simplify the framework without compromising performance. The regression models also demonstrated commendable accuracy, with predictions based on accurate classifier outputs exhibiting a minimum MAPE of 2.3%. This performance showcases the framework’s potential as a reliable predictive component within a digital twin, supporting operational decisions with high confidence. The observed increase in prediction error near the maximum power level, attributed to the logarithmic normalization technique, and the occasional 20% errors in regression models, potentially linked to power state clustering, highlight areas for future refinement. These insights emphasize the need for continuous updates and calibration, consistent with the adaptive nature of digital twins. The insights gleaned from this study lay a robust groundwork for applying deep learning in nuclear reactor modeling and control, paving the way for developing intelligent digital twin systems that promise advancements in operational efficiency and safety.
The continuation of this study will contribute to the development of an intelligent digital twin by integrating advanced deep learning frameworks to enable real-time monitoring and optimization. The proposed digital twin framework aims to not only replicate but also enhance reactor operations through predictive analytics, real-time diagnostics, and adaptive optimization. Future iterations of the proposed framework could incorporate:
@@
  • A robust uncertainty quantification (UQ) framework (derived from authors’ previous studies Kobayashi et al. (2024c); Kumar et al. (2019, 2022)) is essential for ensuring reliable predictions. This could involve Monte Carlo simulations or Bayesian techniques to account for uncertainties in both real-world reactor data and machine learning models.
@@ 5.4 AIO Regressor Model Graph
5.6 over_sample function
1# creates additional samples en masse
2# integral reactivity represented in \$
3def over_sample(dataset, dates, n, change):
4 # create new array for generated data
5 new_data = np.empty(shape=(n,11))
6
7 # reactivities of each core configuration
8 reacs120 = np.array([.03832, .03228, .01778, .00293])/0.006
9 reacs121 = np.array([.03832, .03228, .01778, .00293])/0.006
10 reacs122 = np.array([.03958, .03239, .01778, .00232])/0.006
11 reacs123 = np.array([.0395, .0316, .0181, .0026])/0.006
12 # start dates of each core configuration, in ordinal
13 core121=date(year=2014, month=1, day=15)
14 core121=int(core121.toordinal())
15 core122=date(year=2014, month=10, day=9)
16 core122=int(core122.toordinal())
17 core123=date(year=2014, month=10, day=16)
18 core123=int(core123.toordinal())
19
20 # generate new samples via power ratio
21 for i in range(n):
22 # determine which original sample to use
23 og = np.random.randint(low=0, high=dataset.shape[0])
24 power_i = dataset[og,0]
25 power_f = dataset[og,1]
26 rods_i = dataset[og,2:6]
27 rods_f = dataset[og,6:10]
28 date_code = dataset[og,10]
29 # get the reactivity for this power change
30 day = dates[og]
31 if day<core121:
32 reacs=reacs120
33 config = 0
34 elif day<core122:
35 reacs=reacs121
36 config = 1
37 elif day<core123:
38 reacs=reacs122
39 config = 2
40 else:
41 reacs=reacs123
42 config = 3
43 # generate new initial state
44 new_rods_i, new_power_i, pass_i = power_ratio_sample(rods_i, power_i, reacs, change)
45 # generate new final state
46 new_rods_f, new_power_f, pass_f = power_ratio_sample(rods_f, power_f, reacs, change)
47 # port new information to output array
48 new_data[i,0] = new_power_i # initial power
49 new_data[i,1] = new_power_f # final power
50 new_data[i,2:6] = new_rods_i # initial rod heights
51 new_data[i,6:10] = new_rods_f # final rod heights
52 new_data[i,10] = date_code # ordinal date
53 # if sample doesn’t pass, redo sample
54 if not(pass_i) or not(pass_f):
55 i=i-1
56 return new_data